# RealitySummary: On-Demand Mixed Reality Document Enhancement using Large Language Models

### v1

@@ Abstract.
We introduce RealitySummary, a mixed reality reading assistant that can enhance any printed or digital document using on-demand text extraction, summarization, and augmentation. While augmented reading tools promise to enhance physical reading experiences with overlaid digital content, prior systems have typically required pre-processed documents, which limits their generalizability and real-world use cases. In this paper, we explore on-demand document augmentation by leveraging large language models. To understand generalizable techniques for diverse documents, we first conducted an exploratory design study which identified five categories of document enhancements (summarization, augmentation, navigation, comparison, and extraction). Based on this, we developed a proof-of-concept system that can automatically extract and summarize text using Google Cloud OCR and GPT-4, then embed information around documents using a Microsoft Hololens 2 and Apple Vision Pro. We demonstrate real-time examples of six specific document augmentations: 1) summaries, 2) comparison tables, 3) timelines, 4) keyword lists, 5) summary highlighting, and 6) information cards. Results from a usability study (N=12) and in-the-wild study (N=11) highlight the potential benefits of on-demand MR document enhancement and opportunities for future research.
@@
Mixed Reality; Natural Language Processing; Large Language Models; Augmented Reading
@@
journalyear: 2024copyright: acmlicensedconference: ; May, 2024; $c$s: Human-centered computing Mixed / augmented reality
Figure 1. RealitySummary provides an on-demand reading assistant with embedded mixed reality (MR) text summaries and visual aids. Using markerless tracking, OCR, and GPT-4, the system can provide live summaries, keyword lists, timelines, and text highlighting based on the real-time extraction, analysis, and visualization of any physical or digital document. (Simulated high-resolution illustration based on real Unity assets associated with content in the article shown.)
1. Introduction
“We live in a complex world, filled with myriad objects, tools, toys, and people. [..] Yet, for the most part, our computing takes place sitting in front of, and staring at, a single glowing screen. […] How can we escape from the computer screen and bring these two worlds together?” — Pierre Wellner (Wellner, 1991)
In his pioneering work (Wellner, 1991), Pierre Wellner presented the first demonstration of AR-enhanced documents, arguing that the future of computer interfaces should augment our physical environment, rather than confining us to a flat rectangle screen. This fusion harnessed the dynamic capabilities of digital content while retaining the tangible benefits of physical documents. With recent advances in mixed reality technologies like the Microsoft Hololens and Apple Vision Pro, we are getting closer to realizing his vision where we can naturally augment surrounding paper, books, and documents, as well as content on screens, to enhance our reading and learning experiences (Li et al., 2019; Rajaram and Nebeling, 2022; Chen et al., 2020).
However, most prior systems (Billinghurst et al., 2001; Zhao et al., 2014; Li et al., 2019) have typically required documents to be pre-processed with pre-generated content. The lack of on-demand and real-time content generation for non-prepared documents hinders their generalizability and practicality in real-world applications beyond research labs.
This paper introduces RealitySummary, a mixed reality reading assistant that enhances any printed or digital document using on-demand text extraction, summarization, and augmentation. Our on-demand document augmentation leverages optical character recognition (OCR) and a large language model (LLM) to enhance any given document without prior preparation. To design our system, we first conducted a formative design workshop with six participants to identify common document enhancement techniques across diverse document types. We identified 20 unique techniques across five categories of document enhancement: summarization, augmentation, navigation, comparison, and extraction. Insights from the study led us to incorporate the top six techniques into our system: 1) summaries, 2) comparison tables, 3) timelines, 4) keyword lists, 5) summary highlighting, and 6) information cards.
Based on this insight, we developed our system for the Microsoft Hololens 2 and Apple Vision Pro. The system first captures the document, and then extracts text and the document image using Google’s Cloud OCR API and OpenCV. Given this extracted text, the system generates different augmented content and extracts related images through GPT-4 and Google Image Search. The system also tracks the document based on the extracted document image, rendering the generated content around the document as a spatial 2D canvas. Furthermore, we integrated a speech input using Google’s speech-to-text API, enabling users to activate different summarization features or ask questions related to the current content through voice commands.
To evaluate our system, we conducted two user studies: 1) a usability study with twelve participants, and 2) an in-the-wild study with eleven participants. Both studies highlighted the unique benefits of contextual assistant for diverse set of reading materials without prior preparation. Participants valued the always-on feature for enabling implicit interactions and appreciated the mixed reality interface for facilitating spatial and tangible exploration, significantly enhancing user engagement and the applicability of AI-generated content across various scenarios. Based on the study insights, we also discuss potential risks and future research opportunities of AI-powered mixed reality reading assistants.
Finally, our paper contributes:
  1. (1)
    RealitySummary, a mixed reality reading assistant that uses a large language model to provide on-demand content summarization and augmentation.
  2. (2)
    A set of design features and summarization techniques for mixed reality document enhancement derived from a formative design workshop (N=6).
  3. (3)
    Insights and findings from our usability study (N=12) and in-the-wild study (N=11) that highlight the potential benefits of on-demand mixed reality document enhancements.
@@ 2.1.1. Document Enhancement
However, none of the existing systems have achieved an on-demand document enhancement. Instead, documents required pre-processing with content manually prepared before being embedded in AR. For example, HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document; instead, content was prepared in advance and loaded based on attached fiducial markers. As a result, these systems are not fully adaptable and deployable to real-world scenarios. Some preliminary works have explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) extracting graphs and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information. However, none have yet achieved comprehensive and general-purpose document enhancement. In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models. The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through user evaluations.
2.1.2. Screen-Based Reading Support Tools
@@ 2.2.1. Automatic Summarization
Instead of automatically summarizing the content, the researchers have also explored interactive summarization based on the context and user’s needs. For example, Hoeve et al. (ter Hoeve et al., 2020) studied Q&A interactions to assist users with a document-centric question-answering paradigm, finding that information-seeking was a common query. Semantic Reactor (Google, [n.d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018). VERSE (Vtyurina et al., 2019) is another Q&A system that allows users to ask specific types of questions about a document but is limited by its inability to summarize. Hoeve et al. (ter Hoeve et al., 2020) identified various question types asked by participants regarding documents, with factual/summarization questions being the most frequent. Some question types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions. Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be significantly faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021). Our goal is to build on top of this prior work, developing and exploring automatic and interactive summarization techniques for on-demand and real-time summaries embedded in documents.
3. Exploring Document Enhancement Techniques for Augmented Reading
@@
To explore the possible design of an AI-assisted augmented reading, we conducted formative design workshops with six participants (4 males, 2 females), who had various design backgrounds as undergraduate and graduate HCI researchers.
@@ 3.1. Method
We asked participants to complete a four-part design task where they read and graphically annotated both their own document (Table 1) and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine. After reading the documents, participants brainstormed possible ideas for AI-powered augmentations to enhance their readability. In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then did the same with a printed copy of the Wikipedia article. In the second design task, we gave participants digital copies of the two documents on a virtual whiteboard (Miro) and asked them to design a second set of document enhancements with the assistance of ChatGPT. Each task took approximately 15 minutes, and each session of the entire study took approximately 60 minutes.
Document Types Topic Pages Words
P1 Academic Paper (Horie et al., 2021) HCI 8 2400
P2 Book Chapter (Robinson and Rodrigues, 2022) Religion 6 1800
P3 Academic Paper (Peirson-Smith, 2013) Social Science 4 2000
P4 News Article (NYT, [n.d.]) Non-Fiction 8 1600
P5 Laboratory protocol Engineering 4 1750
P6 Academic Paper (Lin et al., 2018) HCI 4 2200
All Wikipedia Article (wik, [n.d.]) World Affairs 6 4750
Table 1. Documents used by participants in the workshop.
Before commencing the study, we gave participants a verbal walk-through of the study’s purpose and provided a one-page reference containing sample visualizations. We also provided a set of sample prompts for ChatGPT to demonstrate the AI’s summarization and enhancement capabilities. The study followed a talk-aloud protocol and the authors took detailed notes of responses and observations. At the study’s conclusion, the authors collected all sketches authored by the participants.
Upon study completion, all designs elicited from the participants were collated in a single Miro board. Four authors independently categorized and coded those results to identify common themes. We identified twenty distinct types of document enhancements, each with its specific usage, purpose, popularity, and the context in which it was used. Next, all authors discussed the initial design dimensions to refine the consistency and comprehensiveness of the categorization. Finally, four authors collaboratively re-coded the examples to produce a final set of design dimensions.
Figure 2. Document enhancements produced by six participants in our formative design workshop. Highlights Refer to caption indicated the document enhancements implemented in our system.
3.2. Elicited Designs and Findings
We identified twenty unique classes of document enhancements, with each participant generating designs from on average three classes (SD = 1.41). We then identify five high-level categories, which include summarize, compare, augment, extract, and navigate. Figure 2 illustrates the summary of our study results. In this section, we describe each category in more detail.
Category 1. Summarize
One of the most common themes was to expedite the reading process by condensing lengthy text. All participants suggested text-based summaries with different scopes (document-level, page-level, or paragraph-level). Some participants also explored personalized summaries, such as “summarize this for a 5-year-old”, which help understand unfamiliar concepts and fit the unique comprehension levels and interests. Furthermore, many participants recommended the use of text summaries as a reference tool, enabling them to quickly review content from different parts of the document without losing focus on the current page.
Category 2: Compare
The most unique approach was the idea to dynamically change content representation by leveraging AI. For instance, participants envisioned transforming text and numerical data into more visual and spatial formats, such as tables, mind maps, or graphs. These alternative representations enable users to visually compare and analyze elements that would be less apparent in text. Moreover, these representations could be designed to be fluid and adaptive; for instance, tables or graphs could dynamically update, filter, or highlight based on user queries. Another idea was summarizing all of the chronological information into timelines. Such timelines could also automatically update by adding new content as the user reads through the document. This approach represents a unique design for AI-enabled augmentation, which emphasizes reshaping existing content into more comprehensible and comparable formats, offering an alternative method for active reading.
Category 3. Augment
All participants showed interest in enhancing their content with external data, such as maps and images. For instance, one participant envisioned a feature similar to Apple’s keyword lookup for quick reference to unfamiliar words. Such augmented content could come either from external sources (like images and maps), user-generated content (labels and sticky notes), or AI-generated content (converting 2D figures into 3D models). While such ideas have been explored in prior works (Li et al., 2019; Qian et al., 2022), participants also came up with unique AI-driven enhancements beyond simple augmentation. For instance, participants envisioned that by visualizing the importance, relevance, or number of citations for each reference in academic papers, they could quickly determine which paper to read next. Another unique concept involved synthesizing external information with the comparison feature. For example, when looking up the president of Ukraine, participants suggested not just showing a biography image and profile but also visualizing the organizational relationships through a network diagram with other individuals on a page by combining both external information and content on the page.
Category 4. Extract
Participants envisioned interacting with reading materials by extracting elements for persistent references, such as creating keyword lists or extracting figures like scrapbooks. They also considered the possibility of extracting citations from academic articles’ bibliographies or footnotes. For example, one participant envisioned a feature where inline citations could be expanded on paper documents. Furthermore, there was interest in pulling out significant quotes from books over the years, enabling them to accumulate and store these quotes for future use.
Category 5. Navigate
Participants proposed various methods to enhance navigation through dense documents. They suggested the creation of document reading progress indicators to clarify their location within a text. Additionally, there was interest in making a table of contents readily accessible to better comprehend the document’s structure. Another idea involved implementing collapsible headings, allowing users to hide sections they find irrelevant in extensive documents, thereby facilitating a more focused reading experience. One participant suggested making the document searchable in AR. By integrating extraction features, they imagined users verbally dictating a search term and subsequently viewing a list of pages along with a concise summary of that term within the document.
Other Findings Related AI-Driven Augmented Reading
Additional findings include the potential benefits of AI to support idea generation and critical thinking. For example, participants believed that AI-assisted enhancements could deepen understanding by presenting multiple perspectives or providing new insights to enhance critical thinking. On the other hand, participants also identified several possible challenges associated with AI-generated content, including the risk of inaccurate information, instability of output quality, overly lengthy responses, and the possibility of information overload leading to distraction. In summary, while some of the proposed features are beyond the scope of our implementation, these design ideas, along with insights, have informed the development of our system, which we describe next.
4. System and Implementation
Building on insights from our formative study, we developed RealitySummary, a mixed reality reading assistant that utilizes real-time OCR and a large language model.
4.1. System Design
Figure 3 summarizes the overview of our system design. The system first detects and captures the document with the headset’s camera and proceeds to extract text from printed or digital documents using optical character recognition (Google Cloud OCR). Given the extracted text, our system employs the GPT-4 large language model through OpenAI’s API for analyzing and extracting important information from the document. Additionally, we have a middle layer on the server that parses the text using spaCy’s named entity recognition module to extract a list of persons, locations, organizations, and keywords that are used for the web search via Google search API. The processed information is transformed into either a text summary or visual aid with a prepared design template, which we built based on our formative study (Figure 4).
Figure 3. RealitySummary system architecture.
4.1.1. Mixed Reality Headsets and Image Capture
The system then embeds the text and visual output via the mixed reality headset. Since our pipeline is hardware independent, we deployed our system to Microsoft Hololens 2 and Apple Vision Pro. For HoloLens 2, we use an internal camera for image capture, Vuforia for image tracking, and handle content rendering with MRTK3 (Mixed Reality Toolkit 3) SDK [^2] Source:https://github.com/MicrosoftDocs/mixed-reality/blob/docs/mrtk-unity/mrtk3-overview/index.md. For Apple Vision Pro, we implement the same rendering and image tracking features using the RealityKit framework and Swift. Since the Vision Pro does not allow raw camera access, our prototype feeds the camera scene to a MacBook Air via AirPlay, then extracts images via screen capture. Additionally, we offer voice interaction support to modify or create these summary visualizations using Google’s speech-to-text API and a semantic similarity check between the transcript and feature descriptions to detect which feature the user intends to choose.
4.1.2. Text Extraction
We extract text content from physical or digital documents using OCR. We decided to use Google OCR in Google Cloud Vision API for accurate text extraction. OCR output also includes the position of each word, sentence, and paragraph, which helps embed the content for on-document augmentation. However, the OCR processing time is not fast enough for continuous capturing, we only first take a screenshot of the page every 5 seconds, hence, the system only needs to perform OCR once, as long as the page content does not change. We detect whether the user still reads the same page by comparing the new OCR results from the previous one. On top of the repeated capturing, the user can also manually activate the OCR with a voice command.
4.1.3. Document Tracking
We use Vuforia’s markerless image tracking or Vision Pro’s built-in image tracking features to track the document. The captured image is used for both OCR and document image segmentation, which runs in parallel. When the image is captured through the above step, we segment the document boundary with OpenCV rectangle detection. The system then uses the segmented image to generate a dynamic image target in Vuforia and Vision Pro. This allows RealitySummary to track essentially any document with sufficiently distinct visual features without attaching markers in advance. However, the accuracy of the system depends on various factors, such as lighting and tracked image pattern, the user can also manually place and move the content with hands. Recognizing the challenges of image tracking in diverse settings, our Vision Pro implementation also allows viewers to manually anchor content panels to the environment rather than the document. By taking advantage of the platform’s extremely robust spatial tracking, these environmentally-anchored panels (combined with its much higher-resolution displays) dramatically improve the readability of augmented content while reducing instability and jitter caused by moving image targets.
4.1.4. Text Summarization
After text extraction and document localization, the system extracts important information using NLP. We utilize spaCy and the GPT-4 large language model in the server. We use render AR content as a series of 2D canvases. We employ WebSockets communicate between mixed reality headset and the server with JSON format. We also store the OCR results, captured images, and GPT response history in Firebase database. We query multiple prompts to multiple GPT-4 API simultaneously to reduce the generation time. The specific implementation choices for each type of document enhancement are described in the next section.
4.1.5. Speech Input Interaction
The user can enable or disable different summarization panels simply by saying a trigger word like “show me the summary of this page” and “show me the table-of-contents”. Our system leverages Google’s speech-to-text API to convert spoken sentences into text. We leverage spaCy to measure the semantic similarity between the user’s spoken sentence and our predefined list of action descriptions. Beyond a simple summary activation, the user can also ask free-form, open-ended questions about the reading content by simply saying a question, which is directly sent to GPT-4 API along with the OCR results.
4.2. Document Enhancements
In this section, we describe our supported document enhancements. On top of the elicited designs that we derive from our formative study, our design is also inspired by the prior MR- and non-MR-based reading assistant, such as HoloDoc (Li et al., 2019), DuallyNoted (Qian et al., 2022), Marvista (Chen et al., 2022), and SummaryLens (Benharrak et al., 2022). While we acknowledge that individual design elements may not be significantly novel, we believe that our system contributes to the following two aspects: 1) our holistic framework and its evaluation provide a better understanding of document augmentation design space, 2) automated and real-time approach to creating on-demand content opens up a new potential of mixed reality assistant tools.
Each of our document enhancement techniques relies on specific language model queries (more commonly known as prompts, which are documented in our supplemental material), often in combination with additional web services like image search for images and people bio information cards, and mapbox for generating maps. Figure LABEL:fig:documents_user_study shows various reading materials augmented with a variety of information through our system.
Figure 4. RealitySummary document enhancement features. (All images captured live from HoloLens 2.)
Summary
The text summary is extracted directly by using a GPT-4 API call from the Node.js server. We send the entire text of the current page to GPT-4 and get a summary in return, which is then sent to be rendered on HoloLens. We simply prompt GPT-4 asking it to summarize the current text data extracted from the page. This technique is often used in prior summary-based reading assistants like Marvista (Chen et al., 2022). We limit the maximum word length of the summary to less than 150 words, as we learned the user struggles to digest lengthy text through the formative study.
Comparison Table
RealitySummary generates comparison tables based on the user’s prompt. We use GPT-4 to generate a comparison table between the two items along with contextual text from the document similar to generating text (see supplementary material for the prompt). Alternatively, while parsing the document, we prompt GPT-4 beforehand to generate possible comparison tables and store them locally through the session (and query directly from it when comparison action is detected).
Timeline
A timeline is a series of summarised events in a given text. When a user requests a timeline, we prompt GPT-4 to generate a series of events in the form of a timeline based on the user’s query and get a JSON with the relevant data. We then render this using Unity Canvas (Hololens) or HTML Canvas (Vision Pro) in a custom format to make it easy for the user to skim through.
Keyword List
The keyword list is generated using both GPT-4 and spaCy. SpaCy is used for keyword and phrase extraction, which also provides named entity such as date, location, and persons, which will be used for web search. We simply process the text by calling spaCy’s python wrapper using python-shell from our Node.js server and send the result back to the headset for rendering a list of keywords or highlighting them using word/phrase location, similar to keyword-based skimming techniques like ConceptEVA (Zhang et al., 2023).
Summary Highlighting
Summary Highlighting works very similarly to summaries, in that, the user asks to summarize or find more information about a specific phrase, sentence, or paragraph. Once the user wants information about content inside a document, we prompt GPT-4 for a brief summary and use word/phrase, sentence, or paragraph matching to find the location of the content in the document. We then highlight, draw arrows, or annotate based on OCR’s word bounding box results. Other features also leverage word and phrase highlighting to connect relevant information directly to the text in the document.
Information Cards
Inspired by AR-based annotations demonstrated in DigitalDesk (Wellner, 1991) and DuallyNoted (Qian et al., 2022), information cards, as the name suggests, contain information about keywords and people (image with a description), places (a map), and images about common keywords. To reduce the number of requests made to the LLM, we use Google Search API to extract information related to key terms, places, or people. We also use Google Image Search API to source pictures related to terms or phrases on demand. Moreover, we use mapbox API to generate maps. In the rare case that the API fails to provide a response, we query GPT-4 for more information.
4.3. Technical Evaluation
To characterize the technical performance of our prototype, we assessed the reliability of document tracking and overall system latency, as well as the tool’s ability to recognize documents of a variety of different types. To measure the system performances, we also systematically tested RealitySummary with (1) AR tracking, (2) OCR, and (3) summarization relevance/correctness using 20 additional documents (10 physical and 10 on-screen) with greater visual and technical diversity, including posters, newspapers, online articles, patents, magazines, thesis, books, blog posts, etc.
4.3.1. Document Tracking
The AR tracking analysis measured the system’s document tracking reliability with Microsoft Hololens 2 over 60 seconds each. We saw reliable tracking for 14 documents that included graphics (M=92%, SD=5.4% up-time). However, the 6 documents that included only text showed worse performance (M=64%, SD=3.7% up-time) due to their lack of distinct trackable features. AR tracking is not always very reliable, especially with text documents because they have fewer feature points to track. We used system logs to automatically analyze document tracking accuracy throughout the user study. We measured tracking loss whenever the system lost AR rendering or image tracking. We found our system’s tracking to be 72% accurate while tracking documents with only black-and-white text and 92% accurate with tracking documents with some colors and visuals (images, diagrams, charts, etc.). We also found substantial changes in tracking with different lighting conditions (better tracking with better lighting), document image tracking size (large image targets performed much better than smaller documents), and motion (sudden headset movements resulted in temporary tracking loss or flickering).
4.3.2. Applicability of Physical or Digital Content
OCR analysis examined the fraction of words correctly extracted from the documents. Across the 20 documents, we saw extremely high OCR accuracy (M= 97.9%, SD=5.0%). We tested our system’s generalizability and applicability with 20 different document types—newspapers, magazines, academic papers, books, legal documents, and technical documents. Two authors checked each summarization feature on all 20 documents manually to evaluate the applicability and accuracy of the content generated by the large language model. We found our system to be 94% accurate when generalized to other documents in providing correct information based on the context of the document. Summarization analysis verified the relevance and factual correctness of the initial 120 (20 $\times$× 6) summary cards generated by RealitySummary across these 20 documents. Two authors manually examined and verified all documents and cards, and found that 96.77% were both factually correct and relevant.
4.3.3. Latency
In our technical evaluation of latency, we measured the average latency of three different components: AR rendering, network communication, and API responses. Network communication represented our socket.io connection between the Node.js server and the headset client. The system recorded an average latency of 100 ms for rendering AR content in HoloLens and Vision Pro and an average latency of 40 ms for socket.io network connection, which is relatively low and suggests that the rendering and network process is responsive and efficient. On the other hand, API callbacks included two main components, namely external request to the GPT-4 model and pyShell request to our own NLP processing unit built with spaCy. Latency analysis included their API callbacks and wait-time. The average latency for GPT-4 was 2.4 s and NLP Module was 670 ms.
5. Usability Study
We evaluated our system with 12 participants via a think-aloud usability study. Our goal was to qualitatively assess our system’s usability, examine participants’ reactions to our document enhancements, and elicit new opportunities for MR reading assistants.

### v2

# RealitySummary: Exploring On-Demand Mixed Reality Text Summarization and Question Answering using Large Language Models

@@ Abstract.
Large Language Models (LLMs) are gaining popularity as tools for reading and summarization aids. However, little is known about their potential benefits when integrated with mixed reality (MR) interfaces to support everyday reading assistants. We developed RealitySummary, an MR reading assistant that seamlessly integrates LLMs with always-on camera access, OCR-based text extraction, and augmented spatial and visual responses in MR interfaces. Developed iteratively, RealitySummary evolved across three versions, each shaped by user feedback and reflective analysis: 1) a preliminary user study to understand user perceptions (N=12), 2) an in-the-wild deployment to explore real-world usage (N=11), and 3) a diary study to capture insights from real-world work contexts (N=5). Our findings highlight the unique advantages of combining AI and MR, including an always-on implicit assistant, minimal context switching, and spatial affordances, demonstrating significant potential for future LLM-MR interfaces beyond traditional screen-based interactions.
@@
Mixed Reality; Large Language Models; Augmented Reading; In-the-Wild Study; Diary Study
@@
journalyear: 2024copyright: acmlicensedconference: arXiv; September, 2024; $c$s: Human-centered computing Mixed / augmented reality
Figure 1. We developed RealitySummary, an on-demand mixed reality reading assistant, over three iterations and conducted three different studies. Through this evolution, we highlight three key findings: 1) unique benefits of MR and LLMs, 2) emergent use cases of MR reading assistants, and 3) design implications for mixed reality reading assistants.
@@ 1. Related Work
We summarize prior work on reading augmentation and content summarization with a focus on interactive approaches.
@@ 1.1. Reading Augmentation
However, none of the existing systems have achieved an on-demand document enhancement. Instead, documents required pre-processing with content manually prepared before being embedded in AR. For example, HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document; instead, content was prepared in advance and loaded based on attached fiducial markers. As a result, these systems are not fully adaptable and deployable to real-world scenarios. Some preliminary works have explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) extracting graphs and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information. However, none have yet achieved comprehensive and general-purpose document enhancement.
Screen-Based Reading Support Tools
@@ 1.2. Content Summarization
Researchers have also explored interactive summarization approaches based on context and users’ needs. For example, Hoeve et al. (ter Hoeve et al., 2020) studied Q&A interactions to assist users with a document-centric question-answering paradigm, finding that information-seeking was a common query. Semantic Reactor (Google, [n. d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018). VERSE (Vtyurina et al., 2019) is another Q&A system that allows users to ask specific types of questions about a document but is limited by its inability to summarize. Hoeve et al. (ter Hoeve et al., 2020) identified various question types asked by participants regarding documents, with factual/summarization questions being the most frequent. Some question types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions. Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be significantly faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021). Our goal is to build on top of this prior work, developing and exploring automatic and interactive summarization techniques for on-demand and real-time summaries embedded in documents.
The limitations of prior works such as Holodoc (Li et al., 2019), as explained before, prevented the study of long term and in-the-wild use of MR reading assistants. To our knowledge GazePointAR (Lee et al., 2024) is the closent work which explored long-term use of pronoun disambiguation in MR voice assistants. Still, it is unclear how can LLM powered MR assistants are used in everyday life to assist users in various types of reading and knowledge work tasks. We aim to fill this gap. In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models. The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through iterative implementation and multiple user evaluations.
2. Overview of Our Research Approach
Figure 2. Overview of our research approach spanning an evolution of our system across three different versions.
To uncover insights into on-demand reading assistants, we employed a three-phase iterative design process to investigate user interactions and experiences with RealitySummary. Our approach consisted of a usability study on the initial prototype (Study 1: RS-Doc), followed by an in-the-wild study on a second prototype (Study 2: RS-Wild), and concluded with an author-conducted diary study using a third prototype (Study 3: RS-Diary). Each phase was designed to build upon the insights gained from the previous round, allowing for continuous refinement of the design and functionality.
We chose this multi-method, iterative approach to capture a comprehensive understanding of user interactions across various contexts. The initial usability study provided controlled insights into the basic functionality and user experience of RealitySummary. The subsequent in-the-wild study allowed us to observe how users interacted with the improved, more reliable prototype in everyday environments, uncovering real-world usage patterns and challenges. Finally, the diary study offered deep, longitudinal insights into the integration of RealitySummary into daily life, allowing for nuanced observations of subtle interaction patterns and evolving user needs. This progression from controlled to naturalistic settings, coupled with iterative prototyping, allowed us to holistically explore the complex dynamics of AR-AI interactions while addressing emerging design considerations and ethical implications at each stage.
3. Study 0: Elicitation Study
@@
Before implementing our system, we conducted a design elicitation study to explore potential designs for an AI-assisted augmented reading assistant. We held formative design workshops with six participants (4 males and 2 females) who had various design backgrounds as undergraduate and graduate HCI researchers. The study involved two tasks, each lasting approximately 15-20 minutes, with the entire session taking about 60 minutes.
@@ 3.1. Method
We asked participants to complete a four-part design task where they read and graphically annotated both a document of their own and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine. After reading, participants brainstormed ideas for AI-powered augmentations to improve readability. In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then repeated the task with a printed copy of the Wikipedia article. In the second task, participants were provided with digital copies of the two documents on a virtual whiteboard (Miro) and asked to design a second set of document enhancements, this time with assistance from ChatGPT.
Before the study began, we briefed participants, gave them a reference page of sample visualizations, and demonstrated prompts for ChatGPT that showcased its summarization and enhancement capabilities. Using a think-aloud protocol, the authors took detailed notes and collected all participant sketches. These designs were later collated on a Miro board, where four authors independently categorized and coded them, identifying 20 distinct types of document enhancements based on usage, purpose, popularity, and context. The authors then discussed and refined these categories, re-coding the examples to produce a final set of enhancement types  (Figure 3).
Figure 3. In our elicitation study, participants identified a variety of potential summary and augmentation types based on sketching and ChatGPT exercises using diverse text documents.
3.2. Results
We identified five high-level categories: summarize, compare, augment, extract, and navigate. The summarize category focused on condensing text for quick reference, ranging from broad document summaries to personalized summaries for different comprehension levels. Summaries were seen as useful reference tools, allowing users to quickly review content without losing focus. The compare category emphasized transforming textual and numerical data into visual formats like tables, graphs, and timelines, which could dynamically update based on user interactions, making comparisons more intuitive. The augment category explored ways to enhance document content with external or AI-generated data, such as adding maps and images or converting 2D figures into 3D models. Advanced ideas included visualizing the importance of academic citations or integrating external information to provide richer context. The extract category focused on enabling users to pull out and store key elements, such as keywords, figures, or quotes, for future reference. Lastly, the navigate category proposed various methods to improve document navigation, such as progress indicators, accessible tables of contents, and using AR to make documents searchable. These insights, alongside considerations of the potential and challenges of AI-driven tools, informed the development of our system.
4. Study 1 RS-Doc: RealitySummary First Prototype and Usability Study
We implemented the first version of our system on a Microsoft Hololens 2 and conducted a preliminary user evaluation. In this section, we present initial insights from using on-demand, mixed reality document summarization and discuss the associated limitations and design implications.
4.1. Implementation
Figure 4. RS-Doc Hololens 2 architecture.
Figure 5. Examples from our RS-Doc Hololens implementation.
The system design for the HoloLens version involves capturing documents using the HoloLens 2’s internal camera and extracting text with Google Cloud OCR. The extracted text is processed using GPT-4 via OpenAI’s API to identify key information, while spaCy’s named entity recognition (NER) module assists in identifying important entities for further web searches. The system employs Vuforia for markerless image tracking and the Mixed Reality Toolkit 3 (MRTK3) SDK for content rendering. Users interact with the system through voice commands processed via Google’s speech-to-text API, allowing them to modify or query visualizations. The system captures images periodically for OCR, with an option for manual activation, and tracks documents by generating dynamic image targets, ensuring accurate document tracking and augmented content rendering.
Informed by the elicitation study, the system supports several document enhancements aimed at improving the reading experience. These enhancements include text summarization, achieved by sending the current page’s text to GPT-4 for a concise summary, limited to 150 words for easier digestion. Comparison tables are generated on demand based on the document parsing, allowing for quick comparisons within the document. Timelines of events are created by prompting GPT-4 to generate a sequence of summarized events. A keyword list is extracted using GPT-4 and spaCy, with spaCy handling named entities for web searches and keyword highlighting. Summary highlighting allows users to obtain brief summaries or additional information on specific phrases, sentences, or paragraphs, which are then visually highlighted. Information cards provide further details on keywords, people, places, or images, leveraging Google Search, Image APIs, and Mapbox for generating maps. All these features are integrated into the HoloLens environment, offering real-time, interactive content within the mixed reality experience.